# Partial-transpose-guided entanglement classes and minimum noise filtering in many-body Gaussian quantum systems

### v1

@@ I Introduction
Beyond an intriguing feature of nature, the unique correlations of entanglement expressed by quantum particles [1, 2, 3, 4, 5] and quantum fields [6, 7, 8, 9, 10, 11] have been identified as a key resource in quantum information processing [12, 13, 14, 15], inspiring opportunities for exploring previously inaccessible phases of matter and dynamical properties of quantum many-body systems upon their incorporation into computational architectures [16, 17, 18, 19, 15]. For pure quantum states, in which all deviation from purity observed in a reduced density matrix, $\rho_{A}=\text{Tr}_{B}\left(\rho\right)$, arises from entanglement spanning the associated bipartition, entanglement can be directly quantified by the eigenspectrum of $\rho_{A}$ [20, 21, 22, 23, 24, 25]. However, for mixed states, in which entropy is preexisting either due to partial measurements or noisy environments, perspectives of several entanglement measures are required to understand the quantum correlations, e.g., the Entanglement of Formation (EOF) [26], logarithmic negativity [27, 28, 29, 30, 31, 32], and distillable entanglement [26], among others [26, 33, 34, 35, 36], many of which are computationally formidable [37, 38]. The lack of degeneracy between such measures can be appreciated from a physical perspective by noting that classical correlations may obscure access to quantum correlations, and thus entanglement resources required for the creation of a mixed state are lower-bounded by, i.e., can exceed, the entanglement resource of the mixed state itself [39, 40, 41, 42]. Though entanglement in two-body and highly symmetric states is relatively well understood amid this complexity, challenges and open questions arise upon reductions of symmetry even in few-body contexts [43, 44, 45, 46, 47, 48, 49, 50]. Focusing on general many-body bipartite Gaussian mixed states, this manuscript identifies two physically relevant and computable entanglement classes that have clear two-body entanglement structure in the crucial contexts of entanglement detection and state preparation.
@@
Continuous variable quantum states [51, 52, 53, 54, 46, 55, 56] are both experimentally accessible, e.g., Refs. [57, 58, 59, 60, 61, 62, 63, 64], and theoretically valuable, with the Gaussian subset commonly serving as a basis of physically motivated leading-order approximations [65] that offer greater tractability for exploring many-body features of quantum correlations. In the many-body context, inspired in part by the ability to transform pure bipartite Gaussian states into a tensor-product series of $(1_{A}\times 1_{B})$ two-mode entangled pairs with local unitaries [66, 67], several transformations have been devised to investigate properties of Gaussian mixed-state entanglement structure [68, 69, 31, 32, 43, 45]. Among these, Ref. [69] identified a local unitary applicable to disjoint pairs of scalar field vacuum regions that produces a complete entanglement detection process through a tensor-product series of $(1_{A}\times 1_{B})$ entangled mode pairs. In the present work, a deeper understanding of this transformation is developed and leveraged to generalize its performance to all possible many-body bipartite Gaussian states. The transformed state is found to be further conducive to the identification of the entanglement class of states composed entirely of optimal $(1_{A}\times 1_{B})$ resources. These observations reinforce the practical value of complementary perspectives provided by local-unitary-based entanglement rearrangements in the understanding of quantum correlations in many-body mixed states.
@@ II Gaussian entanglement consolidation
The phase space of $n$ continuous variable (CV) modes [51, 52, 53, 54, 46, 55, 56] is spanned by $n$ pairs of position and momentum operators, $\hat{\boldsymbol{r}}\equiv\left(\hat{x}_{1},\hat{p}_{1},...,\hat{x}_{n},\hat{p}_{n}\right)^{T}$, with canonical commutation relations (CCRs) forming a symplectic matrix,
@@
where $\nu_{j}$ is the $j^{\text{th}}$ positive symplectic eigenvalue. For pure states, $\nu_{j}=1$, while for mixed states $\nu_{j}\geq 1$. Methods for calculating $S$ from a CM can be found in, for example, Appendix B.2 or Appendix D of Ref. [78].
@@
For Gaussian states, all information of quantum and classical correlations between modes resides in the CM. One informative quantifier of bipartite entanglement for mixed quantum states is the logarithmic negativity [29, 27, 28], $\mathcal{N}$, which is the sum of negative eigenvalues in the partially transposed (PT) density matrix. In the phase space formalism, this translates to the sum of $n_{-}$ PT-symplectic eigenvalues less than one,
@@ II.1 $\mathcal{N}$-SOL: symplectic orthogonality of local $\mathcal{V}_{\mathcal{N}}$
The notation $\text{GS}[\tilde{S}_{\mathcal{V_{\mathcal{N}}},A}]$ indicates that the GS procedure starts with the $\mathcal{V}_{\mathcal{N}}$ subspace and completes the local symplectic with $2\left(n_{A}-n_{-}\right)$ additional vectors, e.g., from $\mathcal{V}_{{{\cal N}\hskip-3.85002pt/}}$, resulting in symplectic orthonormality, Eq. (2). Details of this GS procedure can be found, for example, in Appendix A of Ref. [69]. Consistent with previous terminology [69], each $(1_{A}\times 1_{B})$ TMSVS space, governed by two vectors in $\mathcal{V_{\mathcal{N}}}$ spanning a degenerate $\tilde{\nu}_{-}$ subspace, will be referred as a core pair, while the complementary space of $\left(n-2n_{-}\right)$ CV modes will be referred to as the halo. Among the post-consolidation product of core pairs, if a negative squeezing pair—for which $\tilde{S}$ row vectors in $\mathcal{V}_{\mathcal{N}}\left(\mathcal{V}_{{{\cal N}\hskip-3.85002pt/}}\right)$ are A-B symmetric(anti-symmetric), as shown in Appendix A—is found, $S_{\pi}$ applies a single-mode phase operation by angle $\pi$ to one of the modes in the core pair. The resulting structure of positive squeezing cores will be utilized throughout subsequent derivations. While the central component of Eq. (14) is consistent with procedures established in Ref. [69] from |$\absolutevalue{i\Omega\tilde{\sigma}}$ |, Eq. (14) generalizes applicability of the procedure to the entire $N$-SOL entanglement class, including when $x$-$p$ mixing is present in the CM  [^3] For the special case of A-B symmetric PT CMs, $\tilde{\sigma}_{A}=\tilde{\sigma}_{B}$, with symmetric off-diagonal block, $\tilde{\sigma}_{AB}=\tilde{\sigma}_{AB}^{T}$, and vanishing $x$-$p$ matrix elements, relevant to physical applications of fields and trapped ions, a local consolidating transformation with $S_{A}=S_{B}$ may also be directly calculated through forthcoming techniques as the S that diagonalizes $\tilde{\sigma}_{A}-\tilde{\sigma}_{AB}$..
@@ II.2 Minimum noise filtering (MNF) and separability
with each filtration, $f$, identifying PSD noise $Y_{f}$. Through the iteration, the set $\left\{\sigma^{\prime}_{c_{f}}\right\}$ of TMSVS pure states can recover $n_{-}$ core pairs associated with the $\mathcal{V_{\mathcal{N}}}$ subspace of the CM, along with possible additional cores due to the subtraction of $Y_{r}$ in the rest of the system  [^4] Note that choosing to identify separability with these additional cores can alter $\mathcal{V_{\mathcal{N}}}$ of the CM.. The iteration stops when the identified $\sigma^{\prime}_{h}$ is NPT without $\mathcal{N}$-SOL structure or PPT, being either completely bound entangled [40] or separable across the A-B bipartition, e.g., distinguished by the separability flow of Ref. [79]. With the separability of Eq. (22) identified through MNF, operations on individual core pairs, $\sigma^{\prime}_{c_{f}}$, do not alter the quantum correlations in the rest of the core-halo system.
@@ IV.1 Application: lattice scalar field vacuum
Figure 2: Entanglement quantified by logarithmic negativity as a function of separation $\tilde{r}$ between two regions ($d=10$) of the infinite one-dimensional free lattice scalar field vacuum in the massless limit ($m=10^{-10}$). Curves show the mixed-state negativity $\mathcal{N}$ and the negativities of underlying pure states identified with field-basis volume measurement [78], $\mathcal{N}_{p}^{\phi}$, or with MNF that creates TMSVS structure for dominant cores arising in the iteration, $\mathcal{N}_{p}^{\text{MNF}}$. Vertical lines indicate transitions between the labeled numbers of cores in the mixed state (solid, gray) and the MNF-identified pure state (dashed, teal). Transitions of the three entanglement classes—$\mathcal{N}$-SOL, $\mathcal{N}$IC, and separable—occur at $\tilde{r}=95$ and $\tilde{r}=98$.
@@
For entangled regions of the lattice scalar field vacuum outside the $\mathcal{N}$IC regime, the MNF pure-state identification provides a new upper bound to $\mathcal{N}_{p}^{\text{min}}$ as a function of $\tilde{r}$, shown by the teal line of Fig. 2. Note that the developed MNF is a direct procedure, such that no optimization process is involved in the production of this upper bound. With $\mathcal{N}_{p}^{\text{MNF}}\sim\mathcal{N}$ at long distances, the $\mathcal{N}$IC entanglement structure is found to provide effective leading-order guidance for the decomposition of a range of $\mathcal{N}$-SOL states surrounding the $\mathcal{N}$IC regime. However, at reduced $\tilde{r}$ where impacts of incompatibility with the $\mathcal{N}$IC entanglement structure built into the MNF accumulate, $\mathcal{N}_{p}^{\text{MNF}}$ becomes comparable to poly/log pure state negativities of previous techniques, e.g., the field-basis volume measurement [78] shown by the black line.
@@ V Summary and Outlook
Beyond future verification of the conjectured exponential decay of the Gaussian Entanglement of Formation with respect to the separation between spacelike regions of the free massless scalar field vacuum, the present manuscript supports several directions of advancing understanding and practical applications of quantum correlations. Expanding upon the current filtration process designed to produce optimal $\mathcal{N}$IC decompositions, alternate filtration strategies may be designed, e.g., governed by the substructure of other entanglement witnesses or the conclusive separability flow known for Gaussian quantum states [79], to identify improved pure-state decompositions of $\mathcal{N}$-SOL states not perturbatively close to the $\mathcal{N}$IC regime, e.g., those appearing between free scalar field vacuum regions in the continuum limit. Though the techniques and calculations of this manuscript have focused on the subset of Gaussian CV states, Gaussian approximations and extensions to the non-Gaussian scope will be important to address the entanglement properties of interacting quantum fields.
@@ Acknowledgements.
We thank D. H. Beck and Ignacio Cirac for discussions of perspective at early stages of this research. Presented calculations utilizing up to 1500 digits of precision were facilitated by the Mathematica 14.1 arbitrary precision libraries [108].
@@ Appendix A Gaussian classical mixing channel for a single core
$\tilde{D}_{TMSVS}=\tilde{S}\tilde{\sigma}_{TMSVS}\tilde{S}^{T}=\frac{1}{2}\begin{pmatrix}I_{2}&I_{2}\\ -I_{2}&I_{2}\\ \end{pmatrix}\tilde{\sigma}_{TMSVS}\begin{pmatrix}I_{2}&-I_{2}\\ I_{2}&I_{2}\\ \end{pmatrix}=\begin{pmatrix}e^{2r}&0&0&0\\ 0&e^{2r}&0&0\\ 0&0&e^{-2r}&0\\ 0&0&0&e^{-2r}\\ \end{pmatrix}\ \ \ .$ (28)
@@
For a CM with PT symplectic transformation given by Eq. (31), there hence exists a decomposition in a Gaussian classical mixing channel specified by Eqs. (29) and (32), where the PSD constraints of $Y_{c_{1}}$ follows from physicality of the CM.
@@ C.2 Scalar field $\mathcal{N}$IC decomposition
$Y_{c_{1}}=\begin{pmatrix}6.65484&0.&6.65484&0.\\ 0.&0.021637&0.&-0.021637\\ 6.65484&0.&6.65484&0.\\ 0.&-0.021637&0.&0.021637\\ \end{pmatrix}\ \ \ .$ (52)

### v2

@@ I Introduction
Beyond an intriguing feature of nature, the unique correlations of entanglement expressed by quantum particles [1, 2, 3, 4, 5] and quantum fields [6, 7, 8, 9, 10, 11] have been identified as a key resource in quantum information processing [12, 13, 14, 15], inspiring opportunities for exploring previously inaccessible phases of matter and dynamical properties of quantum many-body systems upon their incorporation into computational architectures [12, 16, 17, 18, 19]. For pure quantum states, in which all deviation from purity observed in a reduced density matrix, $\rho_{A}=\text{Tr}_{B}\left(\rho\right)$, arises from entanglement spanning the associated bipartition, entanglement can be directly quantified by the eigenspectrum of $\rho_{A}$ [20, 21, 22, 23, 24, 25]. However, for mixed states, in which entropy is preexisting either due to partial measurements or noisy environments, perspectives of several entanglement measures are required to understand the quantum correlations, e.g., the Entanglement of Formation (EOF) [26], logarithmic negativity [27, 28, 29, 30, 31, 32], and distillable entanglement [26], among others [26, 33, 34, 35, 36], many of which are computationally formidable [37, 38]. The lack of degeneracy between such measures can be appreciated from a physical perspective by noting that classical correlations may obscure access to quantum correlations, and thus entanglement resources required for the creation of a mixed state are lower-bounded by, i.e., can exceed, the entanglement resource of the mixed state itself [39, 40, 41, 42]. Though entanglement in two-body and highly symmetric states is relatively well understood amid this complexity, challenges and open questions arise upon reductions of symmetry even in few-body contexts [43, 44, 45, 46, 47, 48, 49, 50]. Focusing on general many-body bipartite Gaussian mixed states, this paper identifies two physically relevant and computable entanglement classes that have clear two-body entanglement structure in the crucial contexts of entanglement detection and state preparation.
@@
Continuous variable (CV) quantum states [47, 51, 52, 53, 54, 55, 52, 56] are both experimentally accessible, e.g., Refs. [57, 58, 59, 60, 61, 62, 63, 64], and theoretically valuable, with the Gaussian subset commonly serving as a basis of physically motivated leading-order approximations [65] that offer greater tractability for exploring many-body features of quantum correlations. In the many-body context, inspired in part by the ability to transform pure bipartite Gaussian states into a tensor-product series of $(1_{A}\times 1_{B})$ two-mode entangled pairs with local unitaries [45, 66], several transformations have been devised to investigate properties of Gaussian mixed-state entanglement structure [67, 68, 29, 30, 44, 46]. Among these, Ref. [68] identified a local unitary applicable to disjoint pairs of scalar field vacuum regions that produces a complete entanglement detection process through a tensor-product series of $(1_{A}\times 1_{B})$ entangled mode pairs. In the present work, a deeper understanding of this transformation is developed and leveraged to generalize its performance to all possible many-body bipartite Gaussian states. The transformed state is found to be further conducive to the identification of the entanglement class of states composed entirely of optimal $(1_{A}\times 1_{B})$ resources. These observations reinforce the practical value of complementary perspectives provided by local-unitary-based entanglement rearrangements in the understanding of quantum correlations in many-body mixed states.
@@ II Gaussian entanglement consolidation
The phase space of $n$ CV modes [47, 51, 52, 53, 54, 55, 52, 56] is spanned by $n$ pairs of position and momentum operators, $\hat{\boldsymbol{r}}\equiv\left(\hat{x}_{1},\hat{p}_{1},...,\hat{x}_{n},\hat{p}_{n}\right)^{T}$, with canonical commutation relations (CCRs) forming a symplectic matrix,
@@
where $\nu_{j}$ is the $j^{\text{th}}$ positive symplectic eigenvalue. For pure states, $\nu_{j}=1$, while for mixed states $\nu_{j}\geq 1$. Methods for calculating $S$ from a CM can be found in, for example, Appendix B.2 of this paper or Appendix D of Ref. [77].
@@
For Gaussian states, all information of quantum and classical correlations between modes resides in the CM. One informative quantifier of bipartite entanglement for mixed quantum states is the logarithmic negativity [27, 31, 32], $\mathcal{N}=\log_{2}||\tilde{\rho}||_{1}$, where $\tilde{\rho}$ is the partially transposed (PT) density matrix and $||\cdot||_{1}$ is the trace norm summing the absolute value of eigenvalues. In the phase space formalism, this translates to the sum of $n_{-}$ PT-symplectic eigenvalues less than one,
@@ II.1 $\mathcal{N}$-SOL: symplectic orthogonality of local $\mathcal{V}_{\mathcal{N}}$
The notation $\text{GS}[\tilde{S}_{\mathcal{V_{\mathcal{N}}},A}]$ indicates that the GS procedure starts with the $\mathcal{V}_{\mathcal{N}}$ subspace and completes the local symplectic with $2\left(n_{A}-n_{-}\right)$ additional vectors, e.g., from $\mathcal{V}_{{{\cal N}\hskip-3.85002pt/}}$, resulting in symplectic orthonormality, Eq. (2). Details of this GS procedure can be found, for example, in Appendix A of Ref. [68]. Consistent with previous terminology [68], each $(1_{A}\times 1_{B})$ TMSVS space, governed by two vectors in $\mathcal{V_{\mathcal{N}}}$ spanning a degenerate $\tilde{\nu}_{-}$ subspace, will be referred as a core pair, while the complementary space of $\left(n-2n_{-}\right)$ CV modes will be referred to as the halo. Among the post-consolidation product of core pairs, if a negative squeezing pair—for which $\tilde{S}$ row vectors in $\mathcal{V}_{\mathcal{N}}\left(\mathcal{V}_{{{\cal N}\hskip-3.85002pt/}}\right)$ are A-B symmetric(anti-symmetric), as shown in Appendix A—is found, $S_{\pi}=-\mathbb{I}_{2}$ applies a single-mode phase operation by angle $\pi$ to one of the modes in the core pair. The resulting structure of positive squeezing cores will be utilized throughout subsequent derivations. While the central component of Eq. (14) is consistent with procedures established in Ref. [68] from |$\absolutevalue{i\Omega\tilde{\sigma}}$ |, Eq. (14) generalizes applicability of the procedure to the entire $N$-SOL entanglement class, including when $x$-$p$ mixing is present in the CM  [^3] For the special case of A-B symmetric PT CMs, $\tilde{\sigma}_{A}=\tilde{\sigma}_{B}$, with symmetric off-diagonal block, $\tilde{\sigma}_{AB}=\tilde{\sigma}_{AB}^{T}$, and vanishing $x$-$p$ matrix elements, relevant to physical applications of fields and trapped ions, a local consolidating transformation with $S_{A}=S_{B}$ may also be directly calculated through forthcoming techniques as the S that diagonalizes $\tilde{\sigma}_{A}-\tilde{\sigma}_{AB}$..
@@ II.2 Minimum noise filtering (MNF) and separability
with each filtration, $f$, identifying PSD noise $Y_{f}$. Through the iteration, the set $\left\{\sigma^{\prime}_{c_{f}}\right\}$ of TMSVS pure states can recover $n_{-}$ core pairs associated with the $\mathcal{V_{\mathcal{N}}}$ subspace of the CM, along with possible additional cores due to the subtraction of $Y_{r}$ in the rest of the system  [^4] Note that choosing to identify separability with these additional cores can alter $\mathcal{V_{\mathcal{N}}}$ of the CM.. The iteration stops when the identified $\sigma^{\prime}_{h}$ is NPT without core structure or PPT, being either completely bound entangled [41] or separable across the A-B bipartition, e.g., distinguished by the separability flow of Ref. [78]. With the separability of Eq. (22) identified through MNF, operations on individual core pairs, $\sigma^{\prime}_{c_{f}}$, do not alter the quantum correlations in the rest of the core-halo system.
@@ IV.1 Application: lattice scalar field vacuum
Figure 2: Entanglement quantified by the logarithmic negativity $\mathcal{N}$ as a function of separation $\tilde{r}$ between two regions ($d=10$) of the infinite one-dimensional free lattice scalar field vacuum in the massless limit ($m=10^{-10}$). Curves show the mixed-state negativity and the negativities of underlying pure states identified with field-basis volume measurement [77], $\mathcal{N}_{p}^{\phi}$, or with MNF that creates TMSVS structure for dominant cores arising in the iteration, $\mathcal{N}_{p}^{\text{MNF}}$. Vertical lines indicate transitions between the labeled numbers of cores in the mixed state (solid, gray) and the MNF-identified pure state (dashed, teal). Transitions of the three entanglement classes—$\mathcal{N}$-SOL, $\mathcal{N}$IC, and separable—occur at $\tilde{r}=95$ and $\tilde{r}=98$.
@@
For entangled regions of the lattice scalar field vacuum outside the $\mathcal{N}$IC regime, the MNF pure-state identification provides a new upper bound to $\mathcal{N}_{p}^{\text{min}}$ as a function of $\tilde{r}$, shown by the teal line of Fig. 2. Note that the developed MNF is a direct procedure, such that no optimization process is involved in the production of this upper bound. With $\mathcal{N}_{p}^{\text{MNF}}\sim\mathcal{N}$ at long distances, the $\mathcal{N}$IC entanglement structure is found to provide effective leading-order guidance for the decomposition of a range of $\mathcal{N}$-SOL states surrounding the $\mathcal{N}$IC regime. However, at reduced $\tilde{r}$ where impacts of incompatibility with the $\mathcal{N}$IC entanglement structure built into the MNF accumulate, $\mathcal{N}_{p}^{\text{MNF}}$ becomes comparable to the polynomially decaying pure-state negativities of previous techniques, e.g., the field-basis volume measurement [77] shown by the black line.
@@ V Summary and Outlook
Beyond future verification of the conjectured exponential decay of the Gaussian Entanglement of Formation with respect to the separation between spacelike regions of the free massless scalar field vacuum, the present paper supports several directions of advancing understanding and practical applications of quantum correlations. Expanding upon the current filtration process designed to produce optimal $\mathcal{N}$IC decompositions, alternate filtration strategies may be designed, e.g., governed by the substructure of other entanglement witnesses or the conclusive separability flow known for Gaussian quantum states [78], to identify improved pure-state decompositions of $\mathcal{N}$-SOL states not perturbatively close to the $\mathcal{N}$IC regime, e.g., those appearing between free scalar field vacuum regions in the continuum limit. Though the techniques and calculations of this paper have focused on the subset of Gaussian CV states, Gaussian approximations and extensions to the non-Gaussian scope will be important to address the entanglement properties of interacting quantum fields.
@@ Acknowledgements.
We thank D. H. Beck and Ignacio Cirac for discussions of perspective at early stages of this research. Interactions with the latter were supported by the Quantum Computing Methods for High Energy Physics program at the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311. Presented calculations utilizing up to 1500 digits of precision were facilitated by the Mathematica 14.1 arbitrary precision libraries [107].
@@ Appendix A Gaussian classical mixing channel for a single core
$\tilde{D}_{TMSVS}=\tilde{S}\tilde{\sigma}_{TMSVS}\tilde{S}^{T}=\frac{1}{2}\begin{pmatrix}\mathbb{I}_{2}&\mathbb{I}_{2}\\ -\mathbb{I}_{2}&\mathbb{I}_{2}\\ \end{pmatrix}\tilde{\sigma}_{TMSVS}\begin{pmatrix}\mathbb{I}_{2}&-\mathbb{I}_{2}\\ \mathbb{I}_{2}&\mathbb{I}_{2}\\ \end{pmatrix}=\begin{pmatrix}e^{2r}&0&0&0\\ 0&e^{2r}&0&0\\ 0&0&e^{-2r}&0\\ 0&0&0&e^{-2r}\\ \end{pmatrix}\ \ \ .$ (28)
@@
For a CM with PT symplectic transformation given by Eq. (31), there hence exists a decomposition in a Gaussian classical mixing channel specified by Eqs. (29) and (32), where the PSD constraints of $Y_{c_{1}}$ follow from physicality of the CM.
@@ C.2 Scalar field $\mathcal{N}$IC decomposition
$Y_{c_{1}}=\begin{pmatrix}6.65484&0&6.65484&0\\ 0&0.021637&0&-0.021637\\ 6.65484&0&6.65484&0\\ 0&-0.021637&0&0.021637\\ \end{pmatrix}\ \ \ .$ (52)